Chemical Reactions: Metathesis and Redox


Part A : Matathesis Reations

1.) 1.0 M sodium carbonate + 6.0 M hydrochloric acid (mix in a small test tube)

Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


2.) 0.1 M sodium carbonate + 0.1 M copper (II) nirate


Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


3.) 0.1 M copper (II) nitrate + 0.1 M sodium phosphate


Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


4.) 0.1 M sodium phosphate + 0.1 M iron (III) nitrate


Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


5.) 0.1 M iron (III) nitrate + 0.1 M sodium hysroxide


Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


6.) 6 M sodium hydroxide + 6 M hydrochloric acid (mix in a small test tube)


Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


7.) 1.0 M sodium bicarbonate (also known as sodium hydrogen carbonate) + 6.0 M hydrochloric acid (mix in a small test tube)


Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


8.) 1.0 M sodium bicarbonate (also known as sodium hydrogen carbonate) + 3.0 M sulfuric acid (mix in a small test tube)


Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


9.) 0.1 M silver nitrate + 0.1 M potassium chloride


Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


10.) 0.1 M potassium chloride + 0.1 M sodium nitrate


Molecular equation:___________________________________________?


Total ionic equation: ___________________________________________?


Net ionic equation:______________________________________________?


Part B:

B1: reactions of metals and hydrochloric acid


Metal: Reaction (Y/N) Observation(s) Molecular equation:

Net ionic equation:

Ca ? ? ? ?

Cu ? ? ? ?

Al ? ? ? ?

Mg ? ? ? ?

Zn ? ? ? ?


B2: reactions of metals and solutions of metal ions


1.) Al with 0.1 M copper (II) chloride

Molecular equation:__________________________________________?

Net ionic equation: ___________________________________________?


2.) Al with 0.1 M magnesium nitrate

Molecular equation:__________________________________________?

Net ionic equation: ___________________________________________?


3.) Al with 0.1 M zinc nitrate

Molecular equation:__________________________________________?

Net ionic equation: ___________________________________________?


4.) copper + 0.1 M magnesium nitrate

Molecular equation:__________________________________________?

Net ionic equation: ___________________________________________?


5.) copper + 0.1 M zinc nitrate

Molecular equation:__________________________________________?

Net ionic equation: ___________________________________________?


6.) magnesium + 0.1 M copper (II) nitrate

Molecular equation:__________________________________________?

Net ionic equation: ___________________________________________?


7.) magnesium metal + 0.1 M zinc nitrate

Molecular equation:__________________________________________?

Net ionic equation: ___________________________________________?


8.) zinc + 0.1 M copper (II) nitrate

Molecular equation:__________________________________________?

Net ionic equation: ___________________________________________?


9.) zinc + 0.1 M magnesium nitrate

Molecular equation:__________________________________________?

Net ionic equation: ___________________________________________?


questions:

1.) based on the observations from part A what would be a reagent to test for the following ions?


a. chloride ion: ____________?

b. carbonate ion:___________?

c. bicarbonate ion:_________?


2.) would the formation of a precipitate by adding silver nitrate guarantee that the precipitating anion is chloride?


3.) how do the results obtain in the reaction of metals with metal ion solutions compare with expected results from the table of activity series given at the end?




1
Expert's answer
2021-03-05T03:43:20-0500

1) Na2Co3+ 2HCl= 2NaCl+H2Co3


Total Ionic reaction =


2na^+ + co3^-2. = 2H^+ + 2Cl^- = 2Na^+ + 2Cl^- + 2H^+. + co3^-2




Ionic reaction=2H+ + Co3-2= H2CO3

2)Na2Co3+ cu(no3)2= 2Nano3+CuCo3


Total Ionic reaction =2na^+ + co3^-2. +Cu^+ + 2No3^- = 2Na^+ + 2Cl^- + Cu+. + 2No3^-


Ionic reaction=Cu^+ + Co3-2= H2CO3


3) Al+ CuCl2= alcl3+ cu


4) Al+ mag(no3)2= alno3+ mg


3) no reaction


4) ZnCo3+ 2HCl= 2ZnCl+H2Co3


Total Ionic reaction =Zn^+2 + co3^-2. = 2H^+ + 2Cl^- = 2Na^+ + 2Cl^- + 2H^+. + co3^-2




Ionic reaction=2H+ + Co3-2= H2CO3